# Modeling metaorder impact with a Non-Markovian Zero Intelligence model

### v1

@@ Abstract
Devising models of the limit order book that realistically reproduce the market response to exogenous trades is extremely challenging and fundamental in order to test trading strategies. We propose a novel explainable model for small tick assets, the Non-Markovian Zero Intelligence, which is a variant of the well-known Zero Intelligence model. The main modification is that the probability of limit orders’ signs (buy/sell) is not constant but is a function of the exponentially weighted mid-price return, representing the past price dynamics, and can be interpreted as the reaction of traders with reservation prices to the price trend. With numerical simulations and analytical arguments, we show that the model predicts a concave price path during a metaorder execution and to a price reversion after the executions ends, as empirically observed. We analyze in-depth the mechanism at the root of the arising concavity, the components which constitute the price impact in our model, and the dependence of the results on the two main parameters, namely the time scale and the strength of the reaction of traders to the price trend.
@@ 3 The Zero Intelligence model and its limitations
The ZI model [8, 9, 10] was designed with the aim of being as analytically tractable as possible. It allows to simulate the LOB by relying on an extremely simple description: limit orders, cancellations and market orders are considered as three independent Poisson processes, with rates that are independent of prices and sizes. Each order has unitary size and bid or ask sides are equally probable, as well as the price levels. Therefore, only four parameters (the rates $\lambda,\mu,\delta$ and the unitary volume $q_{0}$) need to be estimated from empirical data and this can be performed straightforwardly via elementary computations. We adopt the approach illustrated in [3], that we present in Appendix B.
@@ 4.2 The behavior of $\bar{R}_{t}$ and two equivalent descriptions for its computation
From now on, the description under interest is (i) indeed, $\bar{R}_{t}$ should be computed by relying on all market events’ times since a priori the child MOs’ times are not known. Therefore, we intend $\beta_{1}$ for $\beta$ and $\beta_{1}(\Delta+1)$ for $\beta_{2}$.
@@ 4.3 The stationary regime
We showed the behavior of $\bar{R}_{t}$ in the concave regime however, at some point this quantity stabilizes to a stationary value, as represented in the left panel of Fig. 8. Eq. (4) can be rewritten in the following recursive way:
$\begin{split}\bar{R}_{t}&=r_{t}+\gamma\bar{R}_{t-1}\ \ \ t\geq 1\\ \bar{R}_{0}&=0\end{split}$ (6)
where $\gamma=e^{-\beta}<1$ and $r_{t}=m_{t}-m_{t-1}$. Given what we showed in the preceding subsection, we can assume that $\bar{R}_{t}$ is non-decreasing during the execution of a buy metaorder and we can identify the stationary regime as the set of times $\{t\in[t^{*},T]:\ \bar{R}_{t}\simeq\bar{R}_{t-1}\simeq\bar{R}^{*}>0\}$. From Eq. (6), it follows that in this regime also $r_{t}\rightarrow r^{*}$ and the exponentially weighted mid-price return and the mid-price return are approximately proportional:
$\bar{R}^{*}\simeq\frac{r^{*}}{1-\gamma}.$ (7)
@@ 4.4 The mid-price path between two child MOs
We need to consider that, given a state of the order book at time $t$ with mid-price $m=m(t)$ [^6] In this subsection we adopt this notation for the mid-price at time $t$. However, $m(t)$ is analogous to $m_{t}$., it can switch from the value $m$ to one of the values $m^{\prime}=m+\frac{1}{2},m+1,\ldots,m+\frac{s-1}{2}$, where $s=s(t)$ is the spread value at time $t$, because of a buy LO inside the spread at time $t+1$. On the other hand, if at time $t+1$ we have a sell LO inside the spread, the mid-price decreases to $m^{\prime}=m-\frac{1}{2},m-1,\ldots,m-\frac{s-1}{2}$. Sell (buy) MOs and buy (sell) cancellations make the mid-price decrease (increase) by half the first bid (ask) gap size. In the setting of Algorithm 1, the first gap size can assume values between $1$ and $K-2$ where $K$ is the number of tick levels in the grid which represents the LOB. Therefore, since when LOs inside the spread occur, the value that the mid-price can assume depends on the spread, first we need to derive the master equation which describes the evolution of this quantity. Then, we will be able to derive the master equation related to the mid-price.
@@
where $n_{orders}^{bid_{1}/ask_{1}}(t)$ are the numbers of orders at the best bid/ask price level at time $t$. For small-tick assets we can approximate $\mathbb{P}(q_{best}^{ask/bid}(t)=1)\simeq 1\ \forall t$ and therefore, $n_{orders}^{bid_{1}/ask_{1}}(t)\simeq 1\ \forall t$. Additionally, we consider the probability of having the spread difference which coincides with the first gap size as a Kronecker delta:
@@ 5 Conclusion
Several extensions could be interesting to study in future works. In this paper, we focused on the market impact associated with the executions of metaorders with constant speed but undoubtedly, more complex trading strategies could be investigated. Moreover, we consider a specific choice of our indicator of the past price trend $\bar{R}_{t}$, that is the exponentially weighted mid-price return. However, instead of the exponential damping factor, different kernels, as a power-law function, could be employed. Similarly, we define the time-varying probability that a LO is a sell as a sigmoid function of $\bar{R}_{t}$ but, of course, other functional forms could be considered. Finally, similarly to what we do for LOs, also the cancellation signs’ sampling could be performed via a time-varying probability that depends on the past price trend.

### v2

@@ Abstract
Devising models of the limit order book that realistically reproduce the market response to exogenous trades is extremely challenging and fundamental in order to test trading strategies. We propose a novel explainable model for small tick assets, the Non-Markovian Zero Intelligence, which is a variant of the well-known Zero Intelligence model. The main modification is that the probability of limit orders’ signs (buy/sell) is not constant but is a function of the exponentially weighted mid-price return, representing the past price dynamics, and can be interpreted as the reaction of traders with reservation prices to the price trend. With numerical simulations and analytical arguments, we show that the model predicts a concave price path during a metaorder execution and to a price reversion after the execution ends, as empirically observed. We analyze in-depth the mechanism at the root of the arising concavity, the components which constitute the price impact in our model, and the dependence of the results on the two main parameters, namely the time scale and the strength of the reaction of traders to the price trend.
@@ 3 The Zero Intelligence model and its limitations
The ZI model [8, 9, 10] was designed with the aim of being as analytically tractable as possible. It allows to simulate the LOB by relying on an extremely simple description: LOs, cancellations and MOs are considered as three independent Poisson processes, with rates that are independent of prices and sizes. Each order has unitary size and bid or ask sides are equally probable, as well as the price levels. Therefore, only four parameters (the LO rate $\lambda$, the MO rate $\mu$, the cancellation rate $\delta$ and the unitary volume $q_{0}$) need to be estimated from empirical data and this can be performed straightforwardly via elementary computations. We adopt the approach illustrated in [3], that we present in Appendix B.
@@ 4.2 The behavior of $\bar{R}_{t}$ and two equivalent descriptions for its computation
We notice that the economically relevant description is (i): the price trend indicator $\bar{R}_{t}$ should be computed by relying on all market events’ times since a priori the child MOs’ times are not known.
@@ 4.3 The stationary regime
We showed the behavior of $\bar{R}_{t}$ in the concave regime however, at some point this quantity stabilizes to a stationary value, as represented in the left panel of Fig. 8. In the framework of description (ii), Eq. (5) can be rewritten in the following recursive way:
$\begin{split}\bar{R}_{t_{j}+1}^{(2)}&=r_{t_{j}+1}+e^{-\beta_{2}}\bar{R}_{t_{j}-\Delta}^{(2)}\\ \end{split}$ (6)
where $r_{t_{j+1}}=m_{t_{j+1}}-m_{t_{j}-\Delta}$ is the mid-price change due to child MO $j$. Given what we showed in the preceding subsection, the price trend indicator is non-decreasing during the execution of a buy metaorder and we can identify the stationary regime as the set of times $\{t\in[t^{*},T]:\ \bar{R}_{t}\simeq\bar{R}_{t-1}\simeq\bar{R}^{*}>0\}$. From Eq. (6), it follows that in this regime also $r_{t}\rightarrow r^{*}$ and the exponentially weighted mid-price return and the mid-price return are approximately proportional:
$\bar{R}^{*}\simeq\frac{r^{*}}{1-e^{-\beta_{2}}}.$ (7)
@@
From now on, the description under interest is (i), being the economically relevant one, as previously pointed out. Therefore, we intend $\beta_{1}$ for $\beta$ and $\beta_{1}(\Delta+1)$ for $\beta_{2}$.
@@ 4.4 The mid-price path between two child MOs
We need to consider that, given a state of the order book at time $t$ with mid-price $m=m(t)$ [^6] In this subsection we adopt this notation for the mid-price at time $t$. However, $m(t)$ is analogous to $m_{t}$., it can switch from the value $m$ to one of the values $m^{\prime}=m+\frac{1}{2},m+1,\ldots,m+\frac{s-1}{2}$, where $s=s(t)$ is the spread value at time $t$, because of a buy LO inside the spread at time $t+1$. On the other hand, if at time $t+1$ we have a sell LO inside the spread, the mid-price decreases to $m^{\prime}=m-\frac{1}{2},m-1,\ldots,m-\frac{s-1}{2}$. Sell (buy) MOs and buy (sell) cancellations make the mid-price decrease (increase) by half the first bid (ask) gap size. In the setting of Algorithm 1 (displayed in Appendix B), the first gap size can assume values between $1$ and $K-2$ where $K$ is the number of tick levels in the grid which represents the LOB. Therefore, since when LOs inside the spread occur, the value that the mid-price can assume depends on the spread, first we need to derive the master equation which describes the evolution of this quantity. Then, we will be able to derive the master equation related to the mid-price.
@@
where $n_{orders}^{bid_{1}/ask_{1}}(t)$ are the numbers of orders at the best bid/ask price level at time $t$ and $x_{1}^{b/a}(t)$ are the first gaps in the bid/ask side at time $t$. For small-tick assets we can approximate $\mathbb{P}(q_{best}^{ask/bid}(t)=1)\simeq 1\ \forall t$ and therefore, $n_{orders}^{bid_{1}/ask_{1}}(t)\simeq 1\ \forall t$. Additionally, we consider the probability of having the spread difference which coincides with the first gap size as a Kronecker delta:
@@ 5 Conclusion
Several extensions could be interesting to study in future works. In this paper, we focus on the market impact associated with the executions of metaorders with constant speed but undoubtedly, more complex trading strategies could be investigated. Moreover, we consider a specific choice of our indicator of the past price trend $\bar{R}_{t}$, that is the exponentially weighted mid-price return. However, instead of the exponential damping factor, different kernels, as a power-law function, could be employed. Similarly, we define the time-varying probability that a LO is a sell as a sigmoid function of $\bar{R}_{t}$ but, of course, other functional forms could be considered. Finally, similarly to what we do for LOs, also the cancellation signs’ sampling could be performed via a time-varying probability that depends on the past price trend.